# Client-Side Patching against Backdoor Attacks in Federated Learning

### v1

@@ 3.1 Finding Potential Source-Target Class Pairs
In real-world black-box scenarios, the backdoor triggers and source-target class pairs chosen by the attacker are not explicitly known. Consequently, our defense must identify the a of plausible source-target pairs and the trigger patterns associated with them. To address this, we propose using an iterative optimization approach, similar to PGD, to generate a candidate trigger pattern for each source-target class pair.
@@ 3.2 Model Patching
At the end of the FL training process, a patching step is applied to the global model using the triggers identified and optimized by each client during the FL process. This patching mechanism is designed to neutralize backdoor triggers by disrupting the malicious associations learned by the model that enable the backdoor attack.
@@
where, $\ell$ represents a loss function (e.g., the cross-entropy loss). By minimizing $\mathcal{L}_{\text{patch}}(\theta)$, we adjust the parameters of the model $\theta$ to reduce the sensitivity to backdoor triggers while maintaining the performance on clean samples. Therefore, this patching step ensures that the global model $f_{\theta}$ becomes robust to backdoor attacks without introducing additional computational overhead during the standard FL training process.

### v2

@@ 3.1 Finding Potential Source-Target Class Pairs
In real-world black-box scenarios, the backdoor triggers and source-target class pairs chosen by the attacker are not explicitly known. Consequently, our defense must identify the plausible source-target pairs involved in the attack and the trigger patterns associated with them. To address this, we propose using an iterative optimization approach, similar to PGD, to generate a candidate trigger pattern for each source-target class pair.
@@ 3.2 Model Patching
The previous optimization procedure is performed at each iteration of the FL training process. At the end of this process, a patching step is applied to the global model using the triggers identified and optimized by each client. This patching mechanism is designed to neutralize backdoor triggers by disrupting the malicious associations learned by the model that enable the backdoor attack.
@@
where, $\ell$ represents a loss function (e.g., the cross-entropy loss). By minimizing $\mathcal{L}_{\text{patch}}(\theta)$, we adjust the parameters of the model $\theta$ to reduce the sensitivity to backdoor triggers while maintaining the performance on clean samples.
This patching step ensures that the global model $f_{\theta}$ becomes robust to backdoor attacks by fixing the vulnerable spots that are part of backdoor/hidden functionalities. Through the application of this process at the end of the FL training, we reduce computational overhead during the standard training process and ensure that patching information is kept private on the benign nodes, impeding adversaries to guide their attacks towards non-patched areas.